# Probabilistically Tightened Linear Relaxation-based Perturbation Analysis for Neural Network Verification

### v1

@@ 3.2 Extending Wilks’ Probabilistic Guarantees
Building on the results of Lemma 3.6, we can extend the theoretical guarantees of PT-LiRPA to the entire perturbation region $\mathcal{C}$. Specifically, Lemma 3.6 provides a bound on the maximum error in estimating the true minimum (or maximum) of a given intermediate node $z$. By adding this EVT-based error bound to the initial estimates $\overline{l}$ and $\underline{u}$, obtained from the first random sampling of $n$ inputs from $\mathcal{C}$, we derive a probabilistically tight approximation of the reachable set for $z$ that holds for any ${\bm{x}}\in\mathcal{C}$. Notably, following the asymptotic requirements in [19], we can choose the number of upper order statistics as $\nu(n)=\lfloor n^{\xi}\rfloor$, with $\xi\in(0,1)$, which clearly satisfies the conditions $\nu(n)\to\infty$ and $\nu(n)/n\to 0$.
@@
we can compute a linear lower bound function ${\bm{a}}^{T}_{\texttt{PT-LiRPA}}({\bm{x}})+{\bm{c}}_{\texttt{PT-LiRPA}}$ such that with probability at least $1-2mp$ for any ${\bm{x}}\in\mathcal{C}$ satisfies:
@@
We apply the sampling procedure independently for each of the $m$ neurons, estimating the reachable set using $n$ i.i.d. samples drawn from $\mathcal{C}$. By Theorem 3.7, each reachable set is correct with probability at least $1-2p$. Then we use union bound to estimate the probability that all reachable sets are simultaneously correct, yielding a lower bound $1-2mp$. We note that this analysis does not need independence between estimation errors across neurons. In fact, dependencies may arise due to the layered nature of DNNs, which may lead to compounding errors in deeper layers.
@@
$x_{1}$$x_{2}$$z_{1}^{(1)}$$z_{2}^{(1)}$$\hat{z}_{1}^{(1)}$$\hat{z}_{2}^{(1)}$$z_{1}^{(2)}$$z_{2}^{(2)}$$\hat{z}_{1}^{(2)}$$\hat{z}_{2}^{(2)}$$f(x)$222233-3- 311114444ReLUReLU4444222222-2- 21111ReLUReLU22-2- 21111 $[-2,2]$ $[-1,3]$ ${\bm{z}}^{(1)}$ $[-5,7]$ $[-4.96,6.97]$ $[-5.02,6.99]$ $[-10,18]$ $[-9.91,17.98]$ $[-10.00,18.10]$ $[0,18]$ $[0,17.98]$ $[0,18.10]$ $\hat{{\bm{z}}}^{(1)}$ $[0,7]$ $[0,6.97]$ $[0,6.99]$ ${\bm{z}}^{(2)}$ $[-36,28]$ $[-35.96,21.9]$ $[-36.21,22.28]$ $[0,32]$ $[0,19.91]$ $[0,19.92]$ $\hat{{\bm{z}}}^{(2)}$ $[0,28]$ $[0,21.9]$ $[0,22.28]$ $[0,32]$ $[0,19.91]$ $[0,19.92]$ $\bm{[-33.0,18.9]}$ $[-56,32]$ $[-32.82,18.8]$ $[-34.02,18.83]$ [42.,24.3][-42.,24.3][ - 42 . , 24.3 ] $z^{(3)}$
Figure 4: Toy DNN used in this example. Intervals reported in green are the exact output reachable set computed via MIP, in black are the results of the IBP, and in purple the CROWN ones considering the input $[[-2,2],[-1,3]]$. In red are the reachable sets computed using a naive sampling-based approach of $n=10k$ samples. Finally, in blue, the ones computed using a naive sampling-based approach combined with the EVT error estimation.
@@
We start by computing the estimated reachable sets from a sample-based approach in $\mathcal{C}$ using $n=10k$ samples, which for the Proposition 3.2, with $R=0.999$ and considering the number of neurons in the DNN, are sufficient to have a final confidence $\psi\geq 0.99$. We report in Figure 4 highlighted in red the estimated reachable sets obtained from the propagation of $n$ random samples drawn from $[[-2,2],[-1,3]]$. As we can notice, the bounds are slightly tighter than the overestimated ones obtained from the IBP process. Our intuition is thus that from the computation of $\underline{{\bm{D}}}^{(i)},\overline{{\bm{D}}}^{(i)},\underline{{\bm{b}}}^{(i)},\overline{{\bm{b}}}^{(i)}$ using these tightened bounds we can obtain more accurate lower and upper final linear bounds. For the diagonal matrices and bias vectors, we get:
@@
$\underline{{\bm{A}}}^{(2)}$ $=\underline{{\bm{A}}}^{(3)}\underline{{\bm{D}}}^{(2)}{\bm{W}}^{(2)}=[-1.0275,2.5138],$
$\overline{{\bm{A}}}^{(2)}$ $=\overline{{\bm{A}}}^{(3)}\overline{{\bm{D}}}^{(2)}{\bm{W}}^{(2)}=[2,1],$
$\underline{{\bm{A}}}^{(1)}$ $=\underline{{\bm{A}}}^{(2)}\underline{{\bm{D}}}^{(1)}{\bm{W}}^{(1)}=[-1.2,-0.6],$
$\overline{{\bm{A}}}^{(1)}$ $=\overline{{\bm{A}}}^{(2)}\overline{{\bm{D}}}^{(1)}{\bm{W}}^{(1)}=[0.4013,3.7471],$
$\underline{d}$ $=\underline{{\bm{A}}}^{(3)}\underline{{\bm{b}}}^{(2)}+\underline{{\bm{A}}}^{(2)}\underline{{\bm{b}}}^{(1)}=-30.1983,$
$\overline{d}$ $=\overline{{\bm{A}}}^{(3)}\overline{{\bm{b}}}^{(2)}+\overline{{\bm{A}}}^{(2)}\overline{{\bm{b}}}^{(1)}=12.1841,$
@@
$\underline{f}_{\text{CROWN w/ {PT-LiRPA}}}=\min\limits_{{\bm{x}}\in\mathcal{C}}\underline{{\bm{A}}}^{(1)}({\bm{x}})+\underline{d}$ $=-||\underline{{\bm{A}}}^{(1)}||_{1}\cdot\varepsilon+\underline{{\bm{A}}}^{(1)}{\bm{x}}_{0}+\underline{d}$
@@
$\overline{f}_{\text{CROWN w/ {PT-LiRPA}}}=\max\limits_{{\bm{x}}\in\mathcal{C}}\overline{{\bm{A}}}^{(1)}({\bm{x}})+\overline{d}$ $=||\underline{{\bm{A}}}^{(1)}||_{1}\cdot\varepsilon+\overline{{\bm{A}}}^{(1)}{\bm{x}}_{0}+\overline{d}$
@@
As we can notice, these bounds are significantly tighter than the ones computed using CROWN ($[-42,24.3]$). However, the theoretical guarantees provided by Theorem 3.4 allow us to state that these bounds are probabilistically sound only for a fraction $R$ of the perturbation region $\mathcal{C}$, thus resulting in a slightly weaker guarantee w.r.t. the one provided by existing probabilistic approaches. Nonetheless, we believe that if one accepts the assumption underlying this theoretical guarantee, this approach still presents a valuable and computationally efficient tool for computing probabilistically valid linear output bounds. In the following, we show how to practically extend these theoretical guarantees to the whole perturbation region, exploiting Theorem 3.7.
We start again from computing the estimated reachable sets from a sample-based approach in $\mathcal{C}$. For each estimated lower and upper bound, we compute and add the corresponding error using $\frac{Y_{2}-Y_{1}}{(1-p)^{-a_{l}}-1}$ for the lower and $\frac{Y_{n}-Y_{n-1}}{(1-p)^{-a_{u}}-1}$ for the upper bound, respectively. We report in Figure 4 highlighted in blue the new estimated reachable sets obtained from the propagation of $n$ random samples drawn from $[[-2,2],[-1,3]]$ with the addition of the corresponding error. Hence, we speculate that recomputing $\underline{{\bm{D}}}^{(i)},\overline{{\bm{D}}}^{(i)},\underline{{\bm{b}}}^{(i)},\overline{{\bm{b}}}^{(i)}$ using these new estimated reachable sets we can still obtain more accurate lower and upper final linear bounds w.r.t. the LiRPA-based approaches. In fact, we obtain:
@@
$\underline{f}_{\text{PT-LiRPA}}=\min\limits_{{\bm{x}}\in\mathcal{C}}\underline{{\bm{A}}}^{(1)}({\bm{x}})+\underline{d}$ $=-||\underline{{\bm{A}}}^{(1)}||_{1}\cdot\varepsilon+\underline{{\bm{A}}}^{(1)}{\bm{x}}_{0}+\underline{d}$
@@
$\overline{f}_{\text{PT-LiRPA}}=\max\limits_{{\bm{x}}\in\mathcal{C}}\overline{{\bm{A}}}^{(1)}({\bm{x}})+\overline{d}$ $=||\underline{{\bm{A}}}^{(1)}||_{1}\cdot\varepsilon+\overline{{\bm{A}}}^{(1)}{\bm{x}}_{0}+\overline{d}$
@@
A natural question that arises is whether the results of Theorem 3.7 can be used directly to obtain a tight estimation of the output reachable set, without relying on the LiRPA combination. Although this sampling-based method offers a probabilistic estimate that, with high confidence, contains the entire perturbation region, it may still underestimate the true output bounds due to its reliance on a finite number of samples. For example, the MIP result yields bounds of $[-33.0,18.9]$, and as we can notice in Fig. 4 highlighted in blues, the output reachable set only applying a forward computation of Theorem 3.7 still underestimates the exact upper bound $[-34.02,18.83]$. In contrast, since our method integrates a sampling-based approach with any LiRPA method—which inherently provides sound overestimations—the final computed bounds will always be at least as tight as those obtained through estimation based on a finite number of samples and are likely (with a confidence at least $1-2mp$) to produce valid linear bounds, i.e., not discarded by potential adversarial attacks (in fact, we obtain as final result $[-34.4,24.23]$). Additionally, as emphasized in prior work [51], combining forward and backward analysis typically yields tighter bounds compared to using a simple forward bound computation. This observation further motivated our investigation into how tighter reachable sets can enhance the linearization approaches for verification efficiency.
@@ 4 PT-LiRPA framework for Neural Network Verification
Based on the theoretical results of Sec. 3, we now present in Algorithm 1 the PT-LiRPA approach for the verification process. For the sake of clarity and without loss of generality, we present the procedure applied to the parallel BaB as shown for the optimized LiRPA approach proposed in [52].
Algorithm 1 PT-LiRPA on parallel BaB
1:  Input: A DNN $f$ with $m$ neurons, a region $\mathcal{C}$, maximum error in the confidence $p$, sample size $n$, $\xi\in(0,1)$ for $\nu(n)$ and a batch size $t$.
2:  Output: robust/not-robust with a confidence $\geq 1-2pm$.
@@
4:  $\underline{f}_{\mathcal{C}},\overline{f}_{\mathcal{C}}\leftarrow\texttt{LiRPA}(f,\mathcal{C},$interm_bounds))))
5:  $\mathcal{B}\leftarrow\{(\underline{f}_{\mathcal{C}},\overline{f}_{\mathcal{C}})\}$
6:  while $\mathcal{B}\neq\emptyset$ do
7:     $\mathcal{C}_{1},\dots,\mathcal{C}_{t}\leftarrow\texttt{split}(\mathcal{B},m)$
@@
10:     $\mathcal{B}^{\prime}\leftarrow((\underline{f}_{\mathcal{C}_{1}},\overline{f}_{\mathcal{C}_{1}}),\dots,(\underline{f}_{\mathcal{C}_{t}},\overline{f}_{\mathcal{C}_{t}}))$
11:     $\mathcal{B}\leftarrow\mathcal{B}^{\prime}\setminus\texttt{get\_robust\_domain}(\mathcal{B}^{\prime})$
12:     if $\exists\overline{f}_{\mathcal{C}_{i}}<0$ in $\mathcal{B}$ then
13:        return not robust
14:     end if
15:  end while
@@
Given a DNN $f$ with $m$ neurons and a region of interest $\mathcal{C}$, the verification process typically involves a projected gradient descent (PGD) attack [25]. This attack is employed before, after, or during the BaB procedure to search for potential adversarial inputs within the input region under consideration. Similarly, in our approach, we first begin with a PGD attack, and if no adversarial is found, we proceed with the BaB process. We compute the estimated reachable sets using the get_interm_bounds method (line 3), which exploits the results of Theorem 3.7. We then use these bounds in the linear bounds computation on any existing LiRPA approach (line 4), following the computation shown in the toy example of Sec. 3.3. We store the resulting output bounds $\underline{f}$ and $\overline{f}$ for the region $\mathcal{C}$, namely $\underline{f}_{\mathcal{C}}$ and $\overline{f}_{\mathcal{C}}$ in a set $\mathcal{B}$ of unverified regions (lines 5). We then continue the BaB process by splitting (using the split method) the original region from $\mathcal{B}$ into $t$ sub-regions (line 7). Notably, we can perform the parallel selection and splitting into sub-domains using information on unstable ReLU nodes, as shown in [9, 44], or just on the perturbation region $\mathcal{C}_{i}$ [43]. Once we have the new sub-domains, we recompute the estimated reachable sets in parallel and use these bounds for the new computation of the linear lower and upper bounds for each sub-region (lines 8-10). Finally, we update $\mathcal{B}$ with the resulting unverified sub-domains from the procedure get_robust_domain (line 11). The verification process continues until either $\mathcal{B}$ is empty, returning a robust answer, or we find an adversarial configuration, i.e., there is at least a single sub-domain $\mathcal{C}_{i}$ that presents $\overline{f}<0$, thus returning not robust as the answer (lines 12-13). Following Theorem 3.7, if we reach the emptiness of $\mathcal{B}$ thus, no adversarial examples are found during the verification process, we can state that, with a confidence $\geq 1-2mp$, the DNN is robust for the whole perturbation region $\mathcal{C}$.

### v2

@@ 3.2 Extending Wilks’ Probabilistic Guarantees
Building on the results of Lemma 3.6, we can extend the theoretical guarantees of PT-LiRPA to the entire perturbation region $\mathcal{C}$. Specifically, Lemma 3.6 provides a bound on the maximum error in estimating the true minimum (or maximum) of a given intermediate node $z$. By adding this EVT-based error bound to the initial estimates $\overline{l}$ and $\underline{u}$, obtained from the first random sampling of $n$ inputs from $\mathcal{C}$, we derive a probabilistically tight approximation of the reachable set for $z$ that holds for any ${\bm{x}}\in\mathcal{C}$. Notably, following the asymptotic requirements in (haan2006extreme), we can choose the number of upper order statistics as $\nu(n)=\lfloor n^{\xi}\rfloor$, with $\xi\in(0,1)$, which clearly satisfies the conditions $\nu(n)\to\infty$ and $\nu(n)/n\to 0$.
@@
we can compute a linear lower bound function ${\bm{a}}^{T}_{\texttt{PT-LiRPA}}({\bm{x}})+{\bm{c}}_{\texttt{PT-LiRPA}}$ such that with probability at least 12mp1-2mp for any ${\bm{x}}\in\mathcal{C}$ satisfies:
@@
We apply the sampling procedure independently for each of the $m$ neurons, estimating the reachable set using $n$ i.i.d. samples drawn from $\mathcal{C}$. By Theorem 3.7, each reachable set is correct with probability at least 12p1-2p. Then we use union bound to estimate the probability that all reachable sets are simultaneously correct, yielding a lower bound 12mp1-2mp. We note that this analysis does not need independence between estimation errors across neurons. In fact, dependencies may arise due to the layered nature of DNNs, which may lead to compounding errors in deeper layers.
@@
$x_{1}$$x_{2}$$z_{1}^{(1)}$$z_{2}^{(1)}$$\hat{z}_{1}^{(1)}$$\hat{z}_{2}^{(1)}$$z_{1}^{(2)}$$z_{2}^{(2)}$$\hat{z}_{1}^{(2)}$$\hat{z}_{2}^{(2)}$f(x)f(x)223-31144ReLUReLU44222-211ReLUReLU2-211 [2,2][-2,2] [1,3][-1,3] ${\bm{z}}^{(1)}$ [5,7][-5,7] [4.96,6.97][-4.96,6.97] [5.02,6.99][-5.02,6.99] [10,18][-10,18] $\hskip-8.53581pt[-9.91,17.98]$ $\hskip-8.53581pt[-10.00,18.10]$ $[0,18]$ $[0,17.98]$ $[0,18.10]$ $\hat{{\bm{z}}}^{(1)}$ $[0,7]$ $[0,6.97]$ $[0,6.99]$ ${\bm{z}}^{(2)}$ [36,28][-36,28] $\hskip-8.53581pt[-35.96,21.9]$ $\hskip-8.53581pt[-36.21,22.28]$ $[0,32]$ $[0,19.91]$ $[0,19.92]$ $\hat{{\bm{z}}}^{(2)}$ $[0,28]$ $[0,21.9]$ $[0,22.28]$ $[0,32]$ $[0,19.91]$ $[0,19.92]$ $\bm{[-33,18.9]}$ [56,32][-56,32] [32.82,18.8][-32.82,18.8] [34.02,18.83][-34.02,18.83] [42,24.3][-42,24.3] $z^{(3)}$
Figure 5: Toy DNN used in this example. Intervals reported in green are the exact output reachable set computed via MIP, in black are the results of the IBP, and in purple the CROWN ones considering the input [[2,2],[1,3]][[-2,2],[-1,3]]. In red are the reachable sets computed using a naive sampling-based approach of $n=10k$ samples. Finally, in blue, the ones computed using a naive sampling-based approach combined with the EVT error estimation.
@@
We start by computing the estimated reachable sets from a sample-based approach in $\mathcal{C}$ using $n=10k$ samples, which for the Proposition 3.2, with $R=0.999$ and considering the number of neurons in the DNN, are sufficient to have a final confidence $\psi\geq 0.99$. We report in Figure 5 highlighted in red the estimated reachable sets obtained from the propagation of $n$ random samples drawn from [[2,2],[1,3]][[-2,2],[-1,3]]. As we can notice, the bounds are slightly tighter than the overestimated ones obtained from the IBP process. Our intuition is thus that from the computation of $\underline{{\bm{D}}}^{(i)},\overline{{\bm{D}}}^{(i)},\underline{{\bm{b}}}^{(i)},\overline{{\bm{b}}}^{(i)}$ using these tightened bounds we can obtain more accurate lower and upper final linear bounds. For the diagonal matrices and bias vectors, we get:
@@
$\underline{{\bm{A}}}^{(2)}$ $=\underline{{\bm{A}}}^{(3)}\underline{{\bm{D}}}^{(2)}{\bm{W}}^{(2)}=[-1.0275,2.5138],\qquad\overline{{\bm{A}}}^{(2)}=\overline{{\bm{A}}}^{(3)}\overline{{\bm{D}}}^{(2)}{\bm{W}}^{(2)}=[2,1],$
$\underline{{\bm{A}}}^{(1)}$ $=\underline{{\bm{A}}}^{(2)}\underline{{\bm{D}}}^{(1)}{\bm{W}}^{(1)}=[-1.2,-0.6],\qquad\hskip 21.33955pt\overline{{\bm{A}}}^{(1)}=\overline{{\bm{A}}}^{(2)}\overline{{\bm{D}}}^{(1)}{\bm{W}}^{(1)}=[0.4013,3.7471],$
$\underline{d}$ $=\underline{{\bm{A}}}^{(3)}\underline{{\bm{b}}}^{(2)}+\underline{{\bm{A}}}^{(2)}\underline{{\bm{b}}}^{(1)}=-30.1983,\qquad\hskip 24.18483pt\overline{d}=\overline{{\bm{A}}}^{(3)}\overline{{\bm{b}}}^{(2)}+\overline{{\bm{A}}}^{(2)}\overline{{\bm{b}}}^{(1)}=12.1841,$
@@
[…]
@@
[…]
@@
As we can notice, these bounds are significantly tighter than the ones computed using CROWN ([42,24.3][-42,24.3]). However, the theoretical guarantees provided by Theorem 3.4 allow us to state that these bounds are probabilistically sound only for a fraction $R$ of the perturbation region $\mathcal{C}$, thus resulting in a slightly weaker guarantee w.r.t. the one provided by existing probabilistic approaches. Nonetheless, we believe that if one accepts the assumption underlying this theoretical guarantee, this approach still presents a valuable and computationally efficient tool for computing probabilistically valid linear output bounds. In the following, we show how to practically extend these theoretical guarantees to the whole perturbation region, exploiting Theorem 3.7.
We start again from computing the estimated reachable sets from a sample-based approach in $\mathcal{C}$. For each estimated lower and upper bound, we compute and add the corresponding error using $\frac{Y_{2}-Y_{1}}{(1-p)^{-a_{l}}-1}$ for the lower and $\frac{Y_{n}-Y_{n-1}}{(1-p)^{-a_{u}}-1}$ for the upper bound, respectively. We report in Figure 5 highlighted in blue the new estimated reachable sets obtained from the propagation of $n$ random samples drawn from [[2,2],[1,3]][[-2,2],[-1,3]] with the addition of the corresponding error. Hence, we speculate that recomputing $\underline{{\bm{D}}}^{(i)},\overline{{\bm{D}}}^{(i)},\underline{{\bm{b}}}^{(i)},\overline{{\bm{b}}}^{(i)}$ using these new estimated reachable sets we can still obtain more accurate lower and upper final linear bounds w.r.t. the LiRPA-based approaches. In fact, we obtain:
@@
[…]
@@
[…]
@@
A natural question that arises is whether the results of Theorem 3.7 can be used directly to obtain a tight estimation of the output reachable set, without relying on the LiRPA combination. Although this sampling-based method offers a probabilistic estimate that, with high confidence, contains the entire perturbation region, it may still underestimate the true output bounds due to its reliance on a finite number of samples. For example, the MIP result yields bounds of [33.0,18.9][-33.0,18.9], and as we can notice in Fig. 5 highlighted in blues, the output reachable set only applying a forward computation of Theorem 3.7 still underestimates the exact upper bound [34.02,18.83][-34.02,18.83]. In contrast, since our method integrates a sampling-based approach with any LiRPA method—which inherently provides sound overestimations—the final computed bounds will always be at least as tight as those obtained through estimation based on a finite number of samples and are likely (with a confidence at least 12mp1-2mp) to produce valid linear bounds, i.e., not discarded by potential adversarial attacks (in fact, we obtain as final result [34.4,24.23][-34.4,24.23]). Additionally, as emphasized in prior work (autolirpa), combining forward and backward analysis typically yields tighter bounds compared to using a simple forward bound computation. This observation further motivated our investigation into how tighter reachable sets can enhance the linearization approaches for verification efficiency.
@@ 4 PT-LiRPA Framework for Neural Network Verification
Based on the theoretical results of Sec. 3, we now present in Algorithm 3 the PT-LiRPA approach for the verification process. For the sake of clarity and without loss of generality, we present the procedure applied to the parallel BaB as shown for the optimized LiRPA approach proposed in (acrown). Given a DNN $f$ with $m$ neurons and a region of interest $\mathcal{C}$, the verification process typically involves a projected gradient descent (PGD) attack (pgd).
Algorithm 2 PGD_Attackpgd
1:Input Original input ${\bm{x}}_{0}$, neural network $f$, maximum perturbation $\varepsilon$ to create data range $x_{\min},x_{\max}]$, step size $\alpha$, iterations $T$, random_start
2:Output adversarial example ${\bm{x}}_{\text{adv}}$ with $\|{\bm{x}}_{\text{adv}}-{\bm{x}}_{0}\|_{\infty}\leq\varepsilon$ or original input ${\bm{x}}_{0}$
3:if random_start then
4:  ${\bm{x}}\leftarrow{\bm{x}}_{0}+\operatorname{Uniform}(-\varepsilon,\varepsilon)$
5:else
6:  ${\bm{x}}\leftarrow{\bm{x}}_{0}$
@@
8:for $t\in\{1,\ldots,T\}$ do
9:  if $f({\bm{x}}_{\text{adv}})\leq 0$ then
10:   return ${\bm{x}}_{\text{adv}}$
@@
[…]
14:  ${\bm{x}}_{\text{adv}}\leftarrow\operatorname{clip}({\bm{x}}_{0}+\Delta,x_{\min},x_{\max})$
15:return ${\bm{x}}_{0}$
This attack, reported in Alg. 2 for the sake of completeness, can be employed before, after, or during the BaB procedure to search for potential adversarial inputs within the input region under consideration.
Algorithm 3 PT-LiRPA on parallel BaB
1:Input: A DNN $f$ with NN layers and $m$ neurons, an original input ${\bm{x}}_{0}$, a maximum $\varepsilon$ perturbation to create a perturbation region $\mathcal{C}$, maximum error in the confidence $p$, sample size $n$, $\xi\in(0,1)$ for $\nu(n)$ and a batch size $t$.
[…]
3:if PGD_attack($f,\mathcal{C}$) then
4:  return not robust
[…]
8:while $\mathcal{B}\neq\emptyset$ do
9:  $\mathcal{C}_{1},\dots,\mathcal{C}_{t}\leftarrow\texttt{split}(\mathcal{B},t)$
10:  $\textit{interm\_bounds}_{\mathcal{C}_{1},\dots,\mathcal{C}_{t}}\leftarrow\texttt{get\_interm\_bounds}(f,[\mathcal{C}_{1},\dots,\mathcal{C}_{t}],n,p,\xi)$
[…]
12:  $\mathcal{B}^{\prime}\leftarrow(\underline{f}_{\mathcal{C}_{1}},\overline{f}_{\mathcal{C}_{1}}),\dots,(\underline{f}_{\mathcal{C}_{t}},\overline{f}_{\mathcal{C}_{t}})$
[…]
14:   return not robust
15:  $\mathcal{B}\leftarrow\mathcal{B}^{\prime}\setminus\texttt{get\_robust\_domains}(\mathcal{B}^{\prime})$
@@
In detail, we report in our PT-LiRPA algorithm (Alg. 3), a potential employment of PGD during the verification process. Specifically, the attack is performed before and during the BaB, performing a projected gradient descent search in the $L_{\infty}$ ball $\mathcal{C}=\{{\bm{x}}:\|{\bm{x}}-{\bm{x}}_{0}\|_{\infty}\leq\varepsilon\}$ to find an adversarial input ${\bm{x}}_{\text{adv}}$ that makes the scalar model output non-positive, i.e., $f({\bm{x}}_{\text{adv}})\leq 0$. Starting from either the clean original input or a random uniform perturbation in $\varepsilon,\varepsilon]$ (i.e., a random input vector in the $\mathcal{C}$) the method iteratively evaluates the scalar output, computes its gradient with respect to the input, and takes an $L_{\infty}$-constrained descent step using the elementwise sign of the gradient. After each update, the perturbation is projected back onto the $L_{\infty}$ ball and the input is clipped to the valid data range; the procedure stops early if a negative output is obtained and otherwise runs for at most $T$ iterations.
The main hyperparameters are the maximum perturbation $\varepsilon$, the step size $\alpha$ (typically chosen on the order of $\varepsilon/T$), the maximum iterations $T$, and the optional random start; multiple restarts or momentum can be used to increase attack strength. Success provides a concrete counterexample to robustness within the prescribed $L_{\infty}$ radius, while failure is only a heuristic indication and does not constitute a formal certificate of robustness.
Hence, Alg. 3 begins with a PGD attack (lines 3-5), and if no adversarial is found, we proceed with the Branch-and-Bound process. We compute the estimated reachable sets using the get_interm_bounds method (line 6), which exploits the results of Theorem 3.7 and is reported here below in Alg. 4 for clarity.
Algorithm 4 get_interm_bounds
1:interm_bounds $\leftarrow\{\}$
[…]
3:for each intermediate layer do
4:  $^{}\hat{{\bm{l}}},\hat{{\bm{u}}}\leftarrow\{\}$
5:  for each node $z$ in layer nodes do
[…]
7:   $\overline{l},\underline{u}\leftarrow$ $Y_{1}$, $Y_{n}$
8:   $a_{l},a_{u}\leftarrow\frac{\log(\nu)}{\log\left(\frac{Y_{\nu}-Y_{3}}{Y_{3}-Y_{2}}\right)},\frac{\log(\nu)}{\log\left(\frac{Y_{n-2}-Y_{n-\nu}}{Y_{n-1}-Y_{n-2}}\right)}$
[…]
10:   $\hat{{\bm{l}}},\hat{{\bm{u}}}\leftarrow\hat{{\bm{l}}}\cup\hat{l},\hat{{\bm{u}}}\cup\hat{u}$
[…]
12:return interm_bounds
We then use these bounds in the linear bounds computation on any existing LiRPA approach (line 6), following Alg. 1 of Sec. 2.3 and the computation shown in the toy example of Sec. 3.3. We store the resulting output bounds $\underline{f}$ and $\overline{f}$ for the region $\mathcal{C}$, namely $\underline{f}_{\mathcal{C}}$ and $\overline{f}_{\mathcal{C}}$ in a set $\mathcal{B}$ of unverified regions (line 7).
Algorithm 5 get_robust_domains
1:robust_domains $\leftarrow\{\}$
2:for $(\underline{f}_{\mathcal{C}_{i}},\overline{f}_{\mathcal{C}_{i}})\in\mathcal{B}$ do
3:  if $\underline{f}_{\mathcal{C}_{i}}>0$ then
[…]
5:return robust_domains
@@
We then continue the BaB process by splitting (using the split method) the original region from $\mathcal{B}$ into $t$ sub-regions (line 9). Notably, we can perform the parallel selection and splitting into sub-domains using information on unstable ReLU nodes, as shown in (babSplitRelu; bcrown), or just on the perturbation region $\mathcal{C}_{i}$ (reluval). Once we have the new sub-domains, we recompute the estimated reachable sets in parallel and use these bounds for the new computation of the linear lower and upper bounds for each sub-region, and we update $\mathcal{B}$ with the resulting unverified sub-domains (lines 10-12). At each iteration, the process can end either because there is at least a single sub-domain $\mathcal{C}_{i}\in\mathcal{B}$ that presents $\overline{f}<0$, or a PGD attack succeeds, thus returning not robust as the answer (lines 13-16). Otherwise, the process continues updating $\mathcal{B}$ with the unverified domains using the procedure get_robust_domains (line 15) reported in Alg. 5. Following Theorem 3.7, if we reach the emptiness of $\mathcal{B}$, thus no adversarial examples are found during the verification process and all the sub-domains are evaluated as robust, we can state that, with a confidence $\geq 1-2mp$, the DNN is robust for the whole perturbation region $\mathcal{C}$.